OpenStudy (anonymous):

Write the equation of the given line in slope-intercept form. (Points : 1) I have a graph we suppose to use!

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OK, so what you need to do is pick to points on the line

OpenStudy (anonymous):

This is the attanchment with the choices

OpenStudy (anonymous):

and you know the information i am on number 1.

OpenStudy (anonymous):

We have two that are shown already

OpenStudy (anonymous):

Is it 2.4 and -4.1

OpenStudy (anonymous):

Yep, thats the two points

OpenStudy (anonymous):

So how do i write this is is choice C

OpenStudy (anonymous):

So we need an equation of the form y = mx + c where m is the slope and c is the y-intercept.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I need to plug them in

OpenStudy (anonymous):

So now we need to find all the components, that is, m & c

OpenStudy (anonymous):

C is easy, C (the y-intercept) is the point where the line crosses the y axis.

OpenStudy (anonymous):

Do i idividually do it or i try crmming both of them in one equation

OpenStudy (anonymous):

is it y=1/2x + 3

OpenStudy (anonymous):

One at a time, so first we find C. Can you see on the picture where the line crosses through the y axis?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yep, you are correct, it is y = 1/2x + 3

OpenStudy (anonymous):

Woot!

OpenStudy (anonymous):

So its C what i mean by C is the third answer!

OpenStudy (anonymous):

By the way can yiou help me on four more but help me correct them!

OpenStudy (anonymous):

I am on Number 2

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

Y=-2

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

yes i need help correcting

OpenStudy (anonymous):

The second attanchment is the one problem i need help on!

OpenStudy (anonymous):

Can anyone help me?

OpenStudy (anonymous):

I think it is y = -2

OpenStudy (anonymous):

you there

OpenStudy (anonymous):